Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness
Technical integration and coordination
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther
Not logged in
List
|
Find
|
Login
|
Help
Message ID:
42
Entry time:
13 May 2019, 11:24
Author:
Adriano Violante
Location:
General
Severity:
Information
Subject:
Report after dispatch meeting 09.30
SPB/SFX
The gear ratio of the CSS is unclear but this was already known and discussed with SPB/SFX ==> AL is working on it toghter with AE.
FXE
Karabo for Jungfrau detector has been udpated and works fine. For the LPD some issues were reported but SH commented that the problems arise from the fact that the signal FXE was trying to detect is very low, potentially close to the HW limit od the detector
SQS
MPODS power supplies are behaving strangely (1 channel trips, but the power supply does not send any error message). TB and CY to have a look.
The pnCCD is supposed to be used next week. Meeting after the dispatch meeting with AE, CAS, ITDM(!) to coordinate the efforts.
SCS
MID
Mono needs cooling. Issue will be attacked today/tomorrow.
2 crates for the remote contro of vacuum compoennts are required to run in the intrument. CY wants to be present. There are difficulties finding a suitable time slot both for MS and CY. To be addressed offline.
HED
Smaract stages behave strangely. AE support required.
OPT Lasers
Crates will be switched on today. Meeting JE/OPT lasers today at 11 to coordinate the efforts.
XRO
XPD
VAC
DET
Updated pipeline in SPB has been installed today.
SCS DSSC took the first images (without beam)
AE
ITDM
EET
CAS
SRP
TS
AoB
ELOG V3.1.4-7c3fd00